Tohban Report 2013-09-18

From RHESSI Wiki

Jump to: navigation, search

Tohban Reports
Start Date: 11 Sep 2013
End Date: 18 Sep 2013
Tohban: Hazel Bain
Tohban email:
Next Tohban: Lindsay Glesener
List all reports


Solar Activity

Solar activity this week was very low. There was only one C class flare.

Memory Management

The SSR is emptying at the end of the pass set.

Spacecraft Status

Spacecraft status is nominal. The cold plate 1 and 2 are at 117.8 K and 116.2 K respectively.

Data Gaps

There were no data gaps.

Detector issues

The D9 rear segment noise observed last week remained until the 14th of September. The brief dropouts diminished in strength over the start of the week.

Other notes

The hsi_tohban_flare_counter routine is not working suggesting a hiccup in the data processing pipeline. However all data is visible in the browser.

Spacecraft Management

No spin-ups or special commanding this week.

Decimation Normal/Vigorous
HLAT Decimation Rear decimation weight 6, no front decimation
Night time data (fronts) Enabled, during entire eclipse
Night time data (rears) Enabled, during entire eclipse
Require extra passes? No?
Requirement for moving pointer? No
Attenuator operation Normal
Detector problems? Increased noise in Detector 7 continues.
Personal tools